# On the Robustness of Kolmogorov-Arnold Networks: An Adversarial Perspective

### v1

@@ 4.1 Transferability Metric
Let $\mathcal{A}$ denote the attack method, and let $f_{\theta}$ represent the model under attack. The attack can be formalized as:
@@
In our experimental evaluation we assessed the transferability of adversarial examples across multiple attack techniques, discussed in the preceding section. For each attack, a corresponding transferability metric was computed, as defined in Equation 3, quantifying the extent to which adversarial examples generated by a particular attack can deceive the evaluator model under varying conditions.
@@
This metric quantifies the upper bound of the effectiveness of adversarial attacks in terms of their ability to transfer across models, thereby providing a benchmark for evaluating the robustness of $g_{\phi}$ against the most potent adversarial inputs.
4.2 Transferability Assessment
We focused on quantifying the transferability of adversarial examples across all six model configurations of Table 2. Each model acted both as a generator and an evaluator. To ensure a consistent evaluation basis, we selected a subset of $m=5000$ samples from the validation sets of MNIST, KMNIST, and FashionMNIST, which were correctly classified by both models in the transfer pair: model used for attack, model used for prediction. This subset provided a controlled environment to test the adversarial robustness and transferability of attacks.
The transferability of each pair of models was examined using two equations: Equation 3 was used to assess whether two attacks behaved very differently from each other when applied to the same pair of models; Equation 4 was used to identify the most successful attack and to determine the maximum transferability between two models.
The experimental setup aims to explore transferability not only between different types of models, such as MLPs and KANs, but also among models of varying sizes within the KAN category. This dual focus allows us to discern patterns of vulnerability related to both model architecture and size.
4.3 Analysis of Transfer Attacks
In our analysis of the transferability of adversarial attacks, we observed a consistent pattern: the interaction between the model generating the attack and the model evaluating it remained stable across different types of attacks, as shown in Figure 3 (for KMNIST—below we discuss all three datasets).
Figure 3: Comparison of adversarial attack transferability for the KMNIST dataset, using PGD, C&W, MIM, and FGSM attacks. The x-axis represents the model used for generating the attack, the y-axis shows the model used for prediction, and the z-axis indicates the transferability value. Shades of blue are used to represent scenarios where KANs attack MLPs or vice versa, emphasizing cross-model transferability. Shades of red indicate instances where the type of the attack model is the same as the predictor model, highlighting intra-model transferability.
This observation will further help us understand transferability between KANs and MLPs, and how the size of two different KAN models influence transferability. When adversarial images were generated by an MLP model we observed a markedly higher transferability rate, especially when the evaluator was another MLP model. This suggests that MLPs share common vulnerabilities, making them more susceptible to adversarial examples generated by similar architectures. The homogeneity in their architecture likely contributes to this increased transferability, as adversarial perturbations that exploit weaknesses in one MLP tend to be effective against other MLPs.
In contrast—and somewhat unexpectedly—when adversarial examples were generated by a KAN model, there was no significant increase in transferability between KANs and MLPs. In fact, in some cases, there was more transferability between a KAN-attacked model and an MLP evaluator than between KAN-attacked models and KAN evaluators, as evidenced in Table 4.
Table 4: Maximum transferability for FashionMNIST, KMNIST, and MNIST. The value in row $i$, column $j$ represents the maximum transferability between row-$i$ model and column-$j$ model across four attacks—FGSM, PGD, C&W, MIM—calculated using Equation 4. Unlike robustness tables, the values here indicate the percentage of images misclassified. The ‘Average’ row shows the average transferability for the attacking model, with boldface values denoting the lowest transferability, where a lower value indicates better robustness.
(a) FashionMNIST
$\text{KAN}_{\text{small}}$ $\text{KAN}_{\text{medium}}$ $\text{KAN}_{\text{large}}$ $\text{MLP}_{\text{small}}$ $\text{MLP}_{\text{medium}}$ $\text{MLP}_{\text{large}}$
$\text{KAN}_{\text{small}}$ 100.0 40.5 31.7 81.1 65.8 57.2
$\text{KAN}_{\text{medium}}$ 57.1 100.0 29.2 81.3 69.9 58.3
$\text{KAN}_{\text{large}}$ 38.4 28.1 99.9 59.2 45.1 39.4
$\text{MLP}_{\text{small}}$ 59.5 39.9 31.2 99.7 89.7 75.1
$\text{MLP}_{\text{medium}}$ 55.5 39.9 25.7 98.4 99.8 82.2
$\text{MLP}_{\text{large}}$ 48.2 35.8 26.5 90.8 84.8 100.0
Average ($\downarrow$) 59.78 47.37 40.70 85.08 75.85 68.70
(b) KMNIST
$\text{KAN}_{\text{small}}$ $\text{KAN}_{\text{medium}}$ $\text{KAN}_{\text{large}}$ $\text{MLP}_{\text{small}}$ $\text{MLP}_{\text{medium}}$ $\text{MLP}_{\text{large}}$
$\text{KAN}_{\text{small}}$ 100.0 32.8 18.7 70.7 51.8 45.7
$\text{KAN}_{\text{medium}}$ 29.1 100.0 11.7 62.4 42.9 35.7
$\text{KAN}_{\text{large}}$ 20.7 17.5 99.9 46.4 31.2 31.8
$\text{MLP}_{\text{small}}$ 22.0 16.2 7.4 98.7 83.4 50.2
$\text{MLP}_{\text{medium}}$ 13.2 10.7 4.6 89.8 98.8 43.7
$\text{MLP}_{\text{large}}$ 19.3 17.4 10.8 71.5 59.6 99.6
Average ($\downarrow$) 34.05 32.43 25.52 73.25 61.28 51.12
(c) MNIST
$\text{KAN}_{\text{small}}$ $\text{KAN}_{\text{medium}}$ $\text{KAN}_{\text{large}}$ $\text{MLP}_{\text{small}}$ $\text{MLP}_{\text{medium}}$ $\text{MLP}_{\text{large}}$
$\text{KAN}_{\text{small}}$ 100.0 34.6 21.4 81.4 64.3 58.8
$\text{KAN}_{\text{medium}}$ 27.7 100.0 8.4 68.1 54.6 44.7
$\text{KAN}_{\text{large}}$ 22.2 17.2 99.4 49.0 34.8 31.8
$\text{MLP}_{\text{small}}$ 54.8 37.4 17.8 99.7 95.4 78.4
$\text{MLP}_{\text{medium}}$ 30.1 21.6 7.7 95.8 99.6 72.3
$\text{MLP}_{\text{large}}$ 31.9 25.6 14.0 88.3 84.4 99.9
Average ($\downarrow$) 44.45 39.4 28.12 80.38 72.18 64.32
In this table we also observe the “Average” row, which illustrates the average transferability between the attacked model and each of the evaluator models. We can clearly see that across all datasets the adversarial examples generated by the three KAN models successfully fooled the rest of the models far more concinvingly.
5 Concluding Remarks
The introduction of KANs opens new avenues in adversarial machine learning, presenting new challenges and opportunities for enhancing model security. In our analysis, we explored several significant aspects of KANs that have not been extensively studied before.
First, we assessed the inherent robustness of individual KANs against white-box adversarial attacks. Our results indicate that while KANs initially exhibit vulnerability levels similar to those of traditional MLPs, they demonstrate marked improvement in defense as network size increases. This suggests that KANs could be particularly effective in scenarios where larger models are feasible, and robustness is critical.
Secondly, we investigated the transferability of adversarial examples between KANs and MLPs across six different models on MNIST, FashionMNIST, and KMNIST, revealing that cross-model transferability is generally low, which underscores the unique structural and functional characteristics of KANs compared to MLPs. This low transferability suggests that KANs may process information in fundamentally different ways than MLPs. Our comprehensive experiments and analysis underscore the potential of KANs to enhance security in adversarial machine learning.
However, despite these promising findings, it is important to acknowledge that KANs remain fragile in the face of adversarial attacks. Our analysis shows that even as network size increases, KANs are not impervious to adversarial perturbations. This fragility underscores the necessity for continued research to develop more effective strategies for enhancing the robustness of KANs. Understanding and mitigating these vulnerabilities is essential if KANs are to be reliably deployed in security-critical applications.
As this field continues to evolve, further research is needed to fully understand the capabilities and limitations of KANs, paving the way for their practical use when security and reliability are essential.
Appendix A Image Samples of Attacks
Illustration of PGD, C&W, MIM, and FGSM attacks on a single image from the MNIST dataset (Figure 4). We aim to demonstrate how each attack behaves differently for both MLPs and KANs, and how the model sizes influence the behavior of these attacks. Visually, the perturbations observed for MLPs are very different from those found for KANs.
Figure 4: Illustration of PGD, C&W, MIM, and FGSM attacks on a single image from the MNIST dataset, organized into six rows based on model type and size. The first row corresponds to $\text{MLP}_{\text{large}}$, followed by $\text{KAN}_{\text{large}}$, $\text{MLP}_{\text{Medium}}$, $\text{KAN}_{\text{Medium}}$, $\text{MLP}_{\text{Small}}$, and $\text{KAN}_{\text{Small}}$. Each row displays eight columns, where the left four columns show the adversarial images generated by each attack and the right four columns depict the corresponding perturbations.
Appendix B White-Box Attacks
In this appendix we provide a formal mathematical definition of the white-box adversarial model. Additionally, we present comprehensive descriptions of the white-box attacks evaluated in this study.
B.1 Types of White-Box Attacks
B.1.1 Fast Gradient Sign Method (FGSM)
The Fast Gradient Sign Method (FGSM) [16] generates adversarial examples by adding perturbations in the direction of the gradient of the loss function. Specifically, an adversarial example $x_{\text{adv}}$ is computed as follows:
where $\mathcal{L}$ denotes the loss function of the classifier $C$, $\nabla_{x}\mathcal{L}(x,y;\theta)$ represents the gradient of the loss with respect to the input $x$, and $\epsilon$ is a small perturbation. Notably, in Equation 5, only the sign of the gradient is used, with the magnitude of the perturbation controlled by $\epsilon$. It is crucial to highlight that FGSM is a single-step attack; the adversary computes the gradient by backpropagating through the model once—and directly applies this gradient to perturb the input $x$.
B.1.2 Projected Gradient Descent (PGD)
The Projected Gradient Descent (PGD) method [31] is an iterative extension of the FGSM, designed to generate more robust adversarial examples. The PGD attack iteratively perturbs the input $x$ by repeatedly applying gradient updates and projecting the perturbed image back onto an $\epsilon$-ball around the original input. The adversarial example $x_{\text{adv}}$ after $k$ iterations is computed as:
B.1.3 Carlini & Wagner (C&W)
The Carlini & Wagner (C&W) attack [7] is a powerful optimization-based adversarial attack designed to generate adversarial examples by minimizing a specially crafted loss function while ensuring that the perturbation remains imperceptible. When employing the $\ell_{\infty}$ norm, the attack seeks to find a minimal perturbation $\delta$ such that the resulting adversarial example $x_{\text{adv}}=x+\delta$ misleads the classifier. The optimization problem is formulated as follows:
$\min_{\delta}\|\delta\|_{\infty}+c\cdot f(x+\delta),$ (7)
subject to:
$x_{\text{adv}}=\text{clip}(x+\delta,0,1),$
The attack is typically solved through an iterative optimization procedure, where the perturbation $\delta$ is updated in each iteration to reduce the overall objective function.
B.1.4 Momentum Iterative Method (MIM)
The Momentum Iterative Method (MIM) [14] is an enhancement of the basic iterative methods for generating adversarial examples, incorporating momentum to stabilize the gradient updates and avoid poor local maxima. The MIM attack iteratively updates the adversarial example $x_{\text{adv}}$ by accumulating a momentum term, which helps to amplify gradients that consistently point in the same direction across iterations. The adversarial example $x_{\text{adv}}^{k+1}$ after $k$ iterations is computed as:

### v2

@@ B.2.1 Transfer-based Attacks
Let $\mathcal{A}$ denote the white-box adversarial attack method, and let $f_{\theta}$ represent the model under attack. The attack can be formalized as:
@@
B.2.2 Square Attack
The Square Attack (Andriushchenko et al., 2020) is a query-efficient black-box adversarial attack via random search that operates by iteratively applying random perturbations to the input image. The target is to minimize the following loss:
$L(f(x_{\text{adv}}),y)=f_{y}(x_{\text{adv}})-\max_{k\neq y}f_{k}(x_{\text{adv}}),$ (12)
At each iteration, the attack selects a random patch of the image and applies perturbations to it in the form of a square grid. The perturbations are scaled by a predefined $\epsilon$ value, ensuring that they stay within the specified $\ell_{\infty}$ norm constraint. The adversarial example is updated iteratively as follows:
$x_{\text{adv}}^{k+1}=x_{\text{adv}}^{k}+\Delta_{\text{square}},$ (13)
where $\Delta_{\text{square}}$ represents the perturbation applied to the selected patch. The key advantages of the Square Attack are its simplicity and effectiveness, making it a strong baseline for black-box adversarial evaluation. Additionally, its random search mechanism ensures good exploration of the input space, increasing the chances of finding adversarial examples.
B.2.3 Simple Black-Box Attack (SimBA) in DCT Space
The Simple Black-Box Attack (SimBA) (Tu et al., 2019) is a query-efficient method for crafting adversarial examples. Unlike gradient-based attacks, SimBA perturbs the input image in a random manner (e.g., DCT or Fourier space) and evaluates the effect of the perturbation on the model’s confidence score. If the perturbation reduces the confidence of the correct label, it is retained; otherwise, it is discarded.
In the DCT (Discrete Cosine Transform) space, SimBA exploits the frequency-domain representation of images to prioritize perturbations on low-frequency components. This ensures that the adversarial examples remain perceptually similar to the original images while efficiently reducing the model’s confidence in its predictions.
The SimBA attack updates the adversarial example iteratively as follows:
$\ x_{\text{adv}}^{k+1}\ =x_{\text{adv}}^{k}+\Delta_{\text{DCT}},$ (14)
where $\Delta_{\text{DCT}}$ represents the perturbation in the DCT basis. The step size of the perturbation is controlled by the $\epsilon$ parameter.
SimBA’s simplicity and efficiency make it a widely-used black-box attack, particularly in scenarios where query limits are enforced.
B.2.4 Natural Evolution Strategy (NES)
The Natural Evolution Strategy (NES) (Ilyas et al., 2018) is a gradient-free optimization method that estimates the gradient of a loss function by sampling noise vectors and using finite differences to approximate the effect of perturbations. The method is particularly suited for black-box attacks, where access to model gradients is unavailable. It iteratively updates the adversarial example by approximating the gradient of the loss with respect to the input.
For a given input $x$, the adversarial perturbation is computed by estimating the gradient using a set of $n$ random noise vectors $u_{i}\sim\mathcal{N}(0,I)$, sampled from a standard normal distribution. The update rule is as follows:
^$\hat{\nabla}^{k}=\frac{1}{2\sigma n}\sum_{i=1}^{n}\left[\mathcal{L}(x_{\text{adv}}^{k}+\sigma u_{i},y)-\mathcal{L}(x_{\text{adv}}^{k}-\sigma u_{i},y)\right]u_{i}.$ (15)
Using this gradient estimate ^$\hat{\nabla}^{k}$, the adversarial example $x_{\text{adv}}$ is updated iteratively as follows:
$x_{\text{adv}}^{k+1}=\Pi_{\mathcal{B}_{\epsilon}(x)}\left(x_{\text{adv}}^{k}+\alpha\cdot\operatorname{sign}(\hat{\nabla}^{k})\right).$ (16)